# Adaptive Local Combining with Decentralized Decoding for Distributed Massive MIMO

### v2

@@ Abstract
A major bottleneck in uplink distributed massive multiple-input multiple-output networks is the sub-optimal performance of local combining schemes, coupled with high fronthaul load and computational cost inherent in centralized large scale fading decoding (LSFD) architectures. This paper introduces a decentralized decoding architecture that fundamentally breaks from the LSFD, by allowing each access point (AP) to calculate interference-suppressing local weights independently and apply them to its data estimates before transmission. Furthermore, two generalized local zero-forcing (ZF) frameworks, generalized partial full-pilot ZF (G-PFZF) and generalized protected weak PFZF (G-PWPFZF), are introduced, where each AP adaptively and independently determines its combining strategy through a local sum spectral efficiency (SE) optimization that classifies user equipments (UEs) as strong or weak, eliminating the fixed thresholds used in the PFZF and PWPFZF schemes. To enhance scalability, pilot-dependent combining vectors instead of user-dependent ones are introduced and are shared among users with the same pilot. The closed-form SE expressions corresponding to the proposed schemes are derived. Numerical results show that the proposed schemes consistently outperform fixed-threshold counterparts, while the introduction of local weights yields lower overheads and computation costs with lower performance penalty compared to them.
@@ I Introduction
Distributed massive MIMO (D-mMIMO) networks are a promising framework for next-generation wireless systems, as they can enhance weak users’ signal and guarantee uniform service quality across wide coverage areas by enabling numerous distributed access points (APs) to coherently serve user equipments (UEs) [1, 2]. To fully use the advantages of D-mMIMO, while ensuring scalability, distributed signal processing at the APs is essential [3, 4]. A major challenge in this regard is the development of effective uplink combining schemes that can alleviate inter-user interference within practical limitations.
Centralized zero-forcing (ZF) combining, as studied in [5, 6, 7], requires substantial fronthaul signaling and imposes prohibitively high computational costs. Similarly, centralized minimum mean square error (MMSE) combining [8] suffers from both high computation cost and fronthaul overhead, which makes it impractical for large-scale deployments. To alleviate this, partial MMSE combining is proposed in [9, 3, 10], where the interference from only a subset of UEs is considered when designing the combining vector for each UE. Along similar lines, partial centralized approaches have been explored in [11, 12], where joint ZF is applied across a subset of APs based on stronger channels. These partial centralized ZF schemes reduce the fronthaul load compared to fully centralized ZF, but the overhead still scales with the number of UEs, as for each UE, a subset of APs need to send the channel estimates to the CPU, which limits their suitability for ultra-dense network scenarios. These centralized and partial centralized combining schemes follow the fully centralized architecture as shown in Fig. 1, which has very high computational cost and fronthaul overhead, as for all UEs, the channel needs to be estimated at the CPU with the cooperation of all or a subset of APs.
Several distributed combining approaches have been proposed in the literature. Maximum ratio (MR) combining, introduced in [1], is computationally simple and fronthaul-efficient, but its performance is fundamentally limited by its inability to effectively mitigate inter-user interference. To improve upon the MR, a local partial MMSE scheme is proposed in [3, 10], which achieves superior interference suppression and performance. However, a key drawback of this approach is that closed-form expressions for the spectral efficiency (SE) cannot be derived. As a result, resource allocation and system optimization must rely on computationally expensive Monte Carlo simulations, and the lack of tractable expressions makes theoretical performance analysis and gaining insights into the system behavior difficult. In [1, 3], the authors follow a simple decoding architecture as in Fig. 1, thus reducing the overhead apart from local data estimates, yet it suffers from sub-optimality with respect to centralized decoding.
The local combining schemes such as full-pilot ZF (FZF), partial FZF (PFZF), and protected weak PFZF (PWPFZF) are proposed in [13, 14], offering tractable closed-form SE expressions. In these schemes, each AP classifies the UEs into strong and weak groups, generates one combining vector per orthogonal pilot: the ZF for strong UEs and the MR for weak UEs, and assigns scaled versions to individual UEs based on their estimated channel variances [13]. The fundamental limitations of the PFZF and PWPFZF schemes stem from two factors. First, the rigid, threshold-based user grouping may misclassify UEs, leading to suboptimal interference suppression and inefficient use of spatial resources. Second, all APs are constrained to follow the same combining scheme, which prevents each AP from independently adapting its strategy to local channel conditions and interference patterns, further limiting overall system performance.
The large-scale fading decoding (LSFD) as shown in Fig. 1, is employed at the CPU to enhance uplink performance in D-mMIMO networks [8, 15, 13, 14]. In the optimal LSFD approach [13], each AP must transmit both channel and data estimates for all served UEs to the CPU, effectively doubling the fronthaul load while imposing computational costs that scale with the number of users [16]. To alleviate this burden, partial LSFD is proposed in [15], which reduces computational cost and scalability requirements by considering only partial interference in the combining process. However, partial LSFD still requires matrix inversions of the same dimension as optimal LSFD and transmission of the channel estimates to the CPU. As network size and UE density increase, these operations become increasingly demanding with respect to computation and overhead, significantly limiting the scalability and practicality of the LSFD-based architectures in ultra-dense D-mMIMO deployments.
@@
(a) FullyCentralized(b) LSFD(c) SimpleDecoding(d) Proposed DecentralizedDecoding APs CPU All Data Global Combining Global Estimates APs Local Combining CPU Local estimates $\hat{x}_{mt}$Channel estimates LSFD Weights Final Estimates APs Local Combining CPU $\hat{x}_{mt}$ $\hat{x}_{t}=\sum\limits_{m}\hat{x}_{mt}$ APs Local Combining Local Weights $a_{mt}$ CPU $a_{mt}\hat{x}_{mt}$ $\hat{x}_{t}=\sum\limits_{m}a_{mt}\hat{x}_{mt}$
@@
Figure 1: Uplink processing architectures: (a) Fully centralized, (b) LSFD with coordination, (c) Simple decoding, (d) Proposed decentralized decoding.
@@ I-A Contributions
This work addresses some of the above mentioned issues. Our principal contributions are as follows:
Decentralized Decoding Architecture with Local Weighting: We propose a novel uplink processing architecture as shown in Fig. 1, that completely eliminates the need for centralized LSFD and its associated fronthaul overhead. Central to this architecture is the design of interference-suppressing local weights, computed independently by each AP using only its local channel information. These weights are applied to data estimates before transmission, enabling the central processor to operate as a simple aggregator rather than a coordinator. This eliminates the need for sharing any channel information with the CPU. This transition from a coordinated to a truly decentralized architecture is the key enabler for the practical, large-scale deployment.
Adaptive Combining Strategy: We introduce generalized PFZF (G-PFZF) and generalized PWPFZF (G-PWPFZF). Unlike the conventional PFZF and PWPFZF schemes, where all APs must employ the same scheme, the generalized framework allows each AP to independently adapt its combining strategy. Specifically, under G-PFZF an AP may locally switch among PFZF, FZF, or MR; under G-PWPFZF, it may switch among PWPFZF, FZF, or MR, thereby eliminating rigid threshold-based grouping and unlocking per-AP adaptability for improved scalability and performance.
Novel Decentralized Local Optimization for Pilot Partitioning: We also propose novel decentralized optimization framework where each AP independently partitions pilots (and consequently, their associated UEs) into strong or weak groups based on a local sum SE metric. This pilot-level optimization is solved using projected gradient ascent optimization, maintaining low computational cost while maintaining functional equivalence with user-level partitioning.
Closed-Form Spectral Efficiency Expressions: We derive closed-form expressions for the achievable SE of the proposed system. This analysis is essential due to the new statistical properties introduced by pilot-dependent processing and local weighting, and it provides critical insights for system design without the need for Monte Carlo simulation as in MMSE based schemes.
Performance Analysis: Extensive numerical results demonstrate that our decentralized framework and adaptive grouping strategy achieve good SE with reduced computation cost with the state-of-art baselines. The adaptive UE partitioning outperforms traditional fixed-threshold schemes, while the overall design significantly reduces fronthaul overhead and computational costs in comparison to the LSFD framework and fixed threshold partitioning, validating the approach as a practical solution for large-scale deployment.
@@
Organization: The remainder of the paper is organized as follows. Section II describes the system model, along with the channel estimation and data detection procedures. Section III presents the proposed combining schemes and derives their corresponding closed-form SE expressions. Section IV provides a rigorous performance evaluation. Finally, Section V concludes the paper and outlines potential directions for future work.
@@
Notation: Scalars are denoted in italics (e.g., $x$), vectors by bold lowercase letters (e.g., $\mathbf{x}$), and matrices by bold uppercase letters (e.g., $\mathbf{X}$). The transpose and Hermitian transpose are denoted by $(\cdot)^{T}$ and $(\cdot)^{H}$, respectively, while the complex conjugate of a scalar is written as $(\cdot)^{*}$. The sets of real and complex numbers are denoted by $\mathbb{R}$ and $\mathbb{C}$. The expectation operator is denoted by $\mathbb{E}[\cdot]$. The identity matrix of size $N$ is denoted by $\mathbf{I}_{N}$. A complex Gaussian random vector $\mathbf{x}$ with mean $\bm{\mu}$ and covariance matrix $\mathbf{K}$ is denoted as $\mathbf{x}\sim\mathcal{CN}(\bm{\mu},\mathbf{K})$. Also, $_{l}$ represent the $l$-th element of the vector $[.]$.
@@ II System Model
The system consist of $T$ UEs with a single antenna and $M$ APs, where each AP has $A$ antennas. Within the coverage area of interest, there is a uniform distribution of both APs and UEs, and APs work together to service the UEs using the same frequency and time resources. Every AP is linked to the CPU by a fronthaul link, so that the network can operate together and process signals. The wireless channel is represented using a block fading model, characterized by a coherence block of length $L_{c}$ symbols of which $L_{p}$ pilot symbols are designated for pilot training. The small-scale fading is represented by a Rayleigh fading vector $\textbf{h}_{mt}\in\mathbb{C}^{A\times 1}$, and the large-scale fading coefficient (LSFC), encompassing both path-loss and shadowing effects, is denoted by $\beta_{mt}$. The channel vector between AP $m$ and UE $t$, encompassing both small- and large-scale fading, is defined as $\mathbf{g}_{mt}=\beta_{mt}^{1/2}\textbf{h}_{mt}\in\mathbb{C}^{A\times 1}$. We suppose that $\textbf{h}_{mt}\sim\mathcal{CN}(0,\textbf{I}_{A})$, for all $m$ and $t$, is an independent and identically distributed complex Gaussian random vector, thus $\mathbf{g}_{mt}\sim\mathcal{CN}(0,\beta_{mt}\textbf{I}_{A})$.
@@ II-A Channel Estimation
A pilot sequence $\sqrt{L_{p}}\bm{\psi}_{i_{t}}\in\mathbb{C}^{L_{p}\times 1}$ is transmitted by each UE $t$ during the uplink training stage, with $|\bm{\psi}_{i_{t}}|^{2}=1$, where $i_{t}$ is the pilot index of UE $t$. At the $m$-th AP, the received signal is:
@@
where $\textbf{N}_{m}\in\mathbb{C}^{A\times L_{p}}$ represents the additive white Gaussian noise (AWGN) matrix characterized by independent and identically distributed complex Gaussian entries. Additionally, $p^{p}_{t}$ represents the normalized pilot power for UE $t$ during pilot transmission. The MMSE estimate of the true channel vector $\mathbf{g}_{mt}$ is provided by [1]:
$\hat{\textbf{g}}_{mt}=\sqrt{p^{p}_{t}L_{p}}\beta_{mt}\theta^{-1}_{mi_{t}}\textbf{y}^{pilot}_{m}\bm{\psi}_{i_{t}},$
@@
The matrix of estimated channels at AP $m$ is denoted by $\mathbf{\hat{G}}_{m}{=}[\hat{\mathbf{g}}_{m1},\hat{\mathbf{g}}_{m2},\ldots,\hat{\mathbf{g}}_{mT}]{\in}\mathbb{C}^{A\times T}$. Owing to pilot reuse among the UEs, the estimated channel vectors $\hat{\mathbf{g}}_{mt}$ and $\hat{\mathbf{g}}_{mk}$ corresponding to two UEs $t$ and $k$ that share the same pilot sequence become linearly dependent. As a result, the columns of $\mathbf{\hat{G}}_{m}$ exhibit linear dependence, rendering the matrix $\mathbf{\hat{G}}_{m}$ rank-deficient. The full rank matrix is designed by removing the linearly dependent columns and by including only those columns that contain linearly independent channel estimates, representing one vector for each pilot sequence. Thus, the matrix $\mathbf{\bar{G}}_{m}$ is designed as:
@@
where $\bm{\Psi}=[\bm{\psi}_{1},\bm{\psi}_{2},\ldots,\bm{\psi}_{L_{p}}]$. The channel estimate $\hat{\mathbf{g}}_{mt}$ is rewritten as:
$\hat{\mathbf{g}}_{mt}=c_{mt}\mathbf{\bar{G}}_{m}\mathbf{e}_{i_{t}},$
where $\mathbf{e}_{i_{t}}$ is the $i_{t}$-th column of identity matrix $\mathbf{I}_{L_{p}}$.

### v3

@@ Abstract
Efficient uplink processing in distributed massive multiple-input multiple-output (D-mMIMO) systems requires both effective local combining and scalable decoding to significantly mitigate inter-user interference. Recent zero-forcing (ZF)-based combining schemes, such as partial full-pilot ZF (PFZF) and protected weak PFZF (PWPFZF), rely on heuristic threshold-based user grouping that may lead to inefficient utilization of spatial degrees of freedom across access points (APs). To address this limitation, we propose adaptive pilot-aware local combining strategies, generalized PFZF (G-PFZF) and generalized PWPFZF (G-PWPFZF), that dynamically allocate spatial degrees of freedom based on local channel conditions and replace heuristic grouping with a decentralized pilot-level optimization framework. Thus providing substantial performance gains over conventional PFZF and PWPFZF. Further, centralized decoding has recently emerged as a promising technique for interference suppression in D-mMIMO systems. However, it incurs substantial fronthaul overhead and computational costs. We develop a decentralized large-scale fading decoding (d-LSFD) scheme in which each AP computes LSFD weights using only locally available channel statistics. We derive a lower bound on the signal-to-interference-plus-noise ratio that explicitly quantifies the performance gap between the proposed d-LSFD scheme and centralized LSFD (c-LSFD), and identifies conditions under which the proposed decentralized solution approaches the centralized optimum. Numerical results demonstrate that the proposed generalized combining and the d-LSFD scheme together achieve significantly higher sum spectral efficiency in comparison to any combination of existing local combining and decoding schemes, while also substantially reducing the computational cost and fronthaul overhead.
@@ I Introduction
Distributed massive MIMO (D-mMIMO) systems have emerged as a promising architecture for next-generation wireless systems, as they can provide uniform service quality across the network by enabling numerous distributed access points (APs) to coherently serve user equipments (UEs) [1, 2]. However, achieving scalable signal processing in uplink D-mMIMO systems remains challenging because centralized processing architectures require significant fronthaul signaling and computational overhead [3, 4].
Several centralized and partially centralized combining architectures have been proposed to mitigate inter-user interference in D-mMIMO systems by jointly processing signals or channel statistics across multiple APs. [5, 6, 7, 3, 8, 9, 10]. While these approaches can achieve strong interference suppression, they require the central processing unit (CPU) to collect signals or channel statistics from multiple APs, resulting in substantial fronthaul overhead and computational costs, as shown in Fig. 1(a). To reduce such costs, a two-layer decoding architecture has been proposed in [1, 3, 8], where each AP performs local combining while the CPU performs a second-stage decoding. Consequently, the overall performance of D-mMIMO systems critically depends on: the design of efficient local combining schemes at the APs and the design of scalable second-stage decoding techniques at the CPU. Local maximum ratio (MR) combining [1], offers low computational complexity, but suffers from limited interference suppression capability. The local partial minimum mean square error (LP-MMSE) combining scheme [3, 8], achieves better interference suppression and performance. However, it does not provide tractable closed-form spectral efficiency (SE) expressions, making system-level analysis and optimization challenging.
To address these limitations of the MR and LP-MMSE combining, recent local combining schemes such as full-pilot ZF (FZF), partial FZF (PFZF), and protected weak PFZF (PWPFZF)  [11, 14], exploit the structure of pilot reuse to construct combining vectors that partially suppress interference while retaining analytical tractability. In these schemes, each AP classifies the UEs into strong and weak groups using large-scale fading thresholds, allowing some spatial degrees of freedom to be allocated for interference suppression while preserving array gain for weaker users. Although these schemes provide better interference suppression compared to the MR combining, their user grouping decisions rely on heuristic threshold-based rules. As a result, the number of spatial degrees of freedom allocated for interference mitigation is fixed by predetermined thresholds rather than being adapted to the local channel and interference conditions at each AP.
To address such shortcomings of these zero-forcing schemes, we propose adaptive pilot-level combining strategies that dynamically allocate spatial degrees of freedom based on locally observed channel conditions. Specifically, we introduce two generalized combining schemes, referred to as generalized PFZF (G-PFZF) and generalized PWPFZF (G-PWPFZF). Since users sharing the same pilot cannot be fully distinguished by an AP, grouping decisions are naturally performed at the pilot-level rather than at the user level. Here, each AP independently partitions its set of pilots into strong and weak groups by solving a local sum SE optimization problem. Assigning a pilot to the strong group dedicates a spatial degree of freedom (DoF) to mitigate interference via ZF combining, whereas assigning a pilot to the weak group preserves that DoF for signal enhancement through MR combining. This pilot-level formulation allows each AP to adaptively operate across a range of combining strategies from pure MR to FZF. Performance evaluation of the proposed G-PFZF and G-PWPFZF combining schemes establishes that they significantly outperform PFZF and PWPFZF by enabling adaptive pilot-level allocation of spatial degrees of freedom.
In addition to efficient local combining, scalable second-stage decoding is essential for practical uplink processing in D-mMIMO systems. In [1, 3], the authors have proposed a simple decoding technique, as shown in Fig. 1(c) which suffers from suboptimal interference suppression performance. This motivates the use of centralized large-scale fading decoding (c-LSFD) [8, 11], which exploits large-scale channel statistics to mitigate inter-user interference across APs, as shown in Fig. 1(b). While user-centric or scalable architectures reduce fronthaul signaling by limiting the set of serving APs for each UE, centralized LSFD (c-LSFD) designs still require the exchange of channel statistics among these APs.
In the optimal c-LSFD scheme [8, 11], each AP forwards both channel statistics and locally detected signals to the CPU, which computes LSFD weights to suppress inter-user interference. However, this approach requires the exchange of large-scale channel statistics and matrix inversions whose dimensions grow with the number of UEs. Partial LSFD (p-LSFD) has been proposed to reduce the processing complexity by considering only dominant interference terms [8, 12]. Nevertheless, p-LSFD still requires centralized decoding and matrix inversions at the CPU. As the network size and user density increase, centralized operations become increasingly expensive in terms of both fronthaul signaling and computational costs, limiting the scalability of the p-LSFD scheme.
To overcome these scalability limitations, a decentralized decoding scheme, as shown in Fig. 1(d), has been proposed in which each AP locally computes decoding weights [13]. This decentralized heuristic LSFD (h-LSFD) scheme relies on heuristic weight selection and does not provide theoretical guarantees on their performance relative to c-LSFD and p-LSFD. Consequently, it remains unclear whether decentralized decoding with h-LSFD can achieve near c-LSFD performance using only locally computable weights without exchanging global channel statistics.
To address this, we propose a principled d-LSFD scheme where each AP computes decoding weights using only locally available channel statistics through a diagonal approximation of the interference covariance matrix. We establish a theoretical connection between c-LSFD and d-LSFD by deriving an explicit SINR bound that quantifies the performance loss due to decentralization. Furthermore, we show that the effectiveness of the proposed d-LSFD scheme is intrinsically linked to the structure of the local combining scheme. It achieves near-optimal performance for combiners that yield a diagonally dominant large-scale interference covariance matrix (e.g., MR/ZF-based), and larger performance gap for combiners with strong global coupling, such as MMSE.
Extensive numerical results demonstrate that the combination of the proposed G-PFZF and G-PWPFZF combining schemes and the proposed d-LSFD scheme, achieves significantly higher sum SE and per-user SE, and substantially reduced computational cost and fronthaul signaling compared to any combination of existing combining schemes and LSFD schemes. These results highlight the effectiveness of combining adaptive pilot-level interference management with d-LSFD for scalable uplink processing in D-mMIMO systems.
@@
(a) (b) (c) (d) APs All Data CPU Global Combining Global Estimates APs Local Combining Local data estimates ($\hat{x}_{mt}$) and Channel estimates CPU LSFD Weights & Final Estimates APs Local Combining $\hat{x}_{mt}$ CPU $\hat{x}_{t}=\sum\limits_{m}\hat{x}_{mt}$ APs Local Combining Local Weights ($a_{mt}$) $a_{mt}\hat{x}_{mt}$ CPU $\hat{x}_{t}{=}\sum\limits_{m}a_{mt}\hat{x}_{mt}$
@@
Figure 1: Uplink processing architectures: (a) Fully centralized, (b) LSFD with coordination, (c) Simple decoding, (d) Decentralized decoding.
@@ I-A Contributions
We address the twin challenges of designing effective local combining and scalable decoding by developing adaptive pilot-level combining schemes and a decentralized LSFD scheme. In particular, our main contributions are as follows:
Adaptive Combining Schemes: We propose generalized PFZF (G-PFZF) and generalized PWPFZF (G-PWPFZF) schemes that replace heuristic user-level grouping with a pilot-level spatial DoF allocation and provide an explicit mapping between pilot grouping decisions and the number of spatial degrees of freedom used for interference suppression.
Decentralized LSFD Scheme: We propose a decentralized-LSFD (D-LSFD) scheme where each AP computes decoding weights using only locally available channel statistics. We derive locally computable d-LSFD weights and establish an explicit SINR bound that quantifies the performance gap between d-LSFD and c-LSFD. The analysis further reveals that d-LSFD is particularly effective for combining schemes that yield diagonally dominant interference covariance matrices, such as ZF- and MR-type combiners.
Closed-form Expressions: Closed-form SINR and corresponding d-LSFD weights expressions are derived for the proposed combining schemes.
Performance Analysis: Extensive numerical results demonstrate that the combination of the proposed G-PFZF/G-PWPFZF combining and the proposed d-LSFD outperforms any combination of traditional combining schemes with existing LSFD schemes by a significant margin, while significantly reducing the computation cost and fronthaul overhead.
@@
Organization: The remainder of the paper is organized as follows. Section II describes the system model, along with the channel estimation and data detection procedures. Section III presents the proposed combining schemes and derives their corresponding closed-form expressions. Section IV provides a rigorous performance evaluation. Finally, Section V concludes the paper and outlines potential directions for future work.
@@
Notation: Scalars are denoted in italics (e.g., $x$), vectors by bold lowercase letters (e.g., $\mathbf{x}$), and matrices by bold uppercase letters (e.g., $\mathbf{X}$). The transpose and Hermitian transpose are denoted by $(\cdot)^{T}$ and $(\cdot)^{H}$, respectively, while the complex conjugate is written as $(\cdot)^{*}$. The sets of real and complex numbers are denoted by $\mathbb{R}$ and $\mathbb{C}$. The expectation operator is denoted by $\mathbb{E}[\cdot]$. The size $N$ identity matrix is denoted by $\mathbf{I}_{N}$. A complex Gaussian random vector $\mathbf{x}$ with mean $\bm{\mu}$ and covariance matrix $\mathbf{K}$ is denoted as $\mathbf{x}\sim\mathcal{CN}(\bm{\mu},\mathbf{K})$. Also, $_{l}$ represent the $l$-th element of the vector $[.]$.
@@ II System Model
The system consists of $T$ UEs with a single antenna and $M$ APs, where each AP has $A$ antennas. Within the coverage area of interest, there is a uniform distribution of both APs and UEs, and APs serve the UEs using the same frequency and time resources. The wireless channel follows a block fading model with a coherence block of length $L_{c}$ symbols, of which $L_{p}$ symbols are allocated for pilot training. The small-scale fading is represented by a Rayleigh fading vector $\textbf{h}_{mt}{\in}\mathbb{C}^{A\times 1}$, and the large-scale fading coefficient (LSFC), encompassing both path-loss and shadowing effects, is denoted by $\beta_{mt}$. The channel vector between AP $m$ and UE $t$, encompassing both small- and large-scale fading, is defined as $\mathbf{g}_{mt}{=}\beta_{mt}^{1/2}\textbf{h}_{mt}{\in}\mathbb{C}^{A\times 1}$. We assume that $\textbf{h}_{mt}{\sim}\mathcal{CN}(0,\textbf{I}_{A})$, for all $m$ and $t$, is an independent and identically distributed (i.i.d) complex Gaussian random vector, thus $\mathbf{g}_{mt}{\sim}\mathcal{CN}(0,\beta_{mt}\textbf{I}_{A})$. To enhance scalability, we consider user-centric design [8], where $\mathcal{T}_{m}$ represents the set of UEs served by AP $m$ and $\mathcal{M}_{t}$ represents the set of APs that are serving UE $t$.
@@ II-A Channel Estimation
A pilot $\sqrt{L_{p}}\bm{\psi}_{i_{t}}{\in}\mathbb{C}^{L_{p}\times 1}$ is transmitted by each UE $t$ during the uplink training stage, with $|\bm{\psi}_{i_{t}}|^{2}{=}1$, where $i_{t}$ is the pilot index of UE $t$. At the $m$-th AP, the received signal is:
@@
where $\textbf{N}_{m}\in\mathbb{C}^{A\times L_{p}}$ represents the additive white Gaussian noise (AWGN) matrix characterized by i.i.d. complex Gaussian entries. Additionally, $p^{p}_{t}$ represents the normalized pilot power for UE $t$ during pilot transmission. The MMSE estimate of the true channel vector $\mathbf{g}_{mt}$ is provided by [1]:
$\hat{\textbf{g}}_{mt}=\sqrt{p^{p}_{t}L_{p}}\beta_{mt}\theta^{-1}_{mi_{t}}\textbf{y}^{pilot}_{m}\bm{\psi}_{i_{t}},$ (1)
@@
The channel estimates matrix at AP $m$ is denoted by $\mathbf{\hat{G}}_{m}{=}[\hat{\mathbf{g}}_{m1},\hat{\mathbf{g}}_{m2},{\ldots},\hat{\mathbf{g}}_{mT}]{\in}\mathbb{C}^{A\times T}$. Owing to pilot reuse, the columns of $\mathbf{\hat{G}}_{m}$ exhibit linear dependence, rendering the matrix $\mathbf{\hat{G}}_{m}$ rank-deficient. The full rank matrix is designed by removing the linearly dependent columns and by including only those columns that contain linearly independent channel estimates, representing one vector for each pilot sequence. Thus, the matrix $\mathbf{\bar{G}}_{m}$ is designed as:
@@
where $\bm{\Psi}=[\bm{\psi}_{1},\bm{\psi}_{2},\ldots,\bm{\psi}_{L_{p}}]$. Then $\hat{\mathbf{g}}_{mt}=c_{mt}\mathbf{\bar{G}}_{m}\mathbf{e}_{i_{t}}$, where $\mathbf{e}_{i_{t}}$ is the $i_{t}$-th column of identity matrix $\mathbf{I}_{L_{p}}$.